A rectangle has a length 9 meters longer than double the width. The perimeter is 210. Find the dimensions of the rectangle
Answer by macston(5194)

W=width; L=length=2W+9m; P=perimeter=2(L+W)=210m
P=2(L+W)
210m=2(L+W) Divide each side by 2.
105m=L+W Substitute for L.
105m=2W+9m+W Subtract 9m from each side.
96m=3W Divide each side by 3.
32m=W ANSWER 1: The width is 32 meters.
L=2W+9m=2(32m)+9m=64m+9m=73m ANSWER 2: The length is 73 meters.
CHECK:
P=2(L+W)
210 m=2(73 m+32 m)
210 m=2(105 m)
210 m=210 m